# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ IV.4 Quantum many-body probes with long-range interactions
Since local interactions are useful in quantum enhanced sensing, it is natural to ask whether criticality in models with longer-range interactions are also beneficial for metrology tasks Monika et al. (2023). Unfortunately, exact results are rare for such systems and approximate methods, such as Density Matrix Renormalization Group (DMRG), are often unsatisfactory for generic Hamiltonians with long-range interactions because of an area law violation. However, QFI has indeed been studied in many systems showing interesting physics. Perhaps the most studied model is the Lipshin-Meshkov-Glick (LMG) Hamiltonian for N sites in 1D, which generalizes the anisotropic XY Hamiltonian by dropping the nearest-neighbour assumption. The Hamiltonian can be written as
$H_{P}=\frac{J}{N}\sum_{i,j,i<j}\left(\sigma_{i}^{x}\sigma_{j}^{x}+\gamma\sigma_{i}^{y}\sigma_{j}^{y}\right)+h\sum_{i}\sigma_{i}^{z},$ (50)
where $\gamma$ is the anisotropy parameter, $J$ is exchange coupling, and $h$ is the magnetic field. This model has been solved through Bethe ansatz (Pan and Draayer, 1999; Morita et al., 2006), or alternatively by a Holstein-Primakoff transformation (Dusuel and Vidal, 2004, 2005). Considering $J$ as the unit of interaction, this model has three estimable parameters, the magnetic field strength $h$, anisotropy $\gamma$, and the temperature $T$ when system is in thermal equilibrium. At zero temperature, this system shows quantum criticality at $h{=}h_{c}{=}J$. In Ref. (Kwok et al., 2008), it has been shown that the QFI with respect to $h$ diverges at criticality as $I_{Q}\sim(h-h_{c})^{-2}$ (namely $\alpha{=}2$) in the symmetric phase ($h/J\geq 1$) and $I_{Q}\sim(h_{c}-h)^{-1/2}$ (namely $\alpha{=}1/2$) in the broken phase ($0\leq h/J\leq 1$). Remarkably, these exponents for QFI do not depend on the anisotropy $\gamma$. While at the critical point $h_{c}$, one obtains $I_{Q}(h_{c})\sim N^{4/3}$, indicating quantum-enhanced sensitivity, two-site reduced density matrices results in $I_{Q}(h_{c})\sim N^{2/3}$ (Ma et al., 2008). In (Salvatori et al., 2014), the authors considered the estimation of $\gamma$ and $T$. In the thermodynamic limit near criticality, the leading contribution to QFI for both scales as $I_{Q}(h_{c})\sim T^{2}$ at both phases, making this system attractive for thermometry. However, QFI for estimating anisotropy $\gamma$ in the thermodynamic limit shows a divergence at $\gamma=1$ and scales differently in the ordered phase ($h>1$) compared to the symmetry-broken phase ($10\leq h<$ 1) following the formula
@@
Garbe et al (Garbe et al., 2022) have recently considered the influence of various driving protocols on QFI-based sensing in LMG and other fully connected systems. In LMG and other fully connected models, all interactions are equally strong, in contrast to models where all interactions except the nearest or next nearest are ignored. Thus, one may wonder what happens for models where long-range interactions are present, but the interaction strength is not distance agnostic, e.g., it decays following a power law (Fernández-Lorenzo et al., 2018). Such models become extremely relevant for ion-trap based platforms, where the decay exponent can be fixed by tuning the lasers Monroe et al. (2021). In Ref. (Yang et al., 2022a), the authors considered a long-range Kitaev chain, which is an extension of the $N$-site tight-binding model, and whose Hamiltonian is given by
$H$ $=-\frac{J}{2}\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j+1}+c_{j+1}^{\dagger}c_{j})-\mu\sum_{j=1}^{N}(c_{j}^{\dagger}c_{j}-\frac{1}{2})$
$+\frac{\Delta}{2}\sum_{j=1}^{N-1}\sum_{l=1}^{N-j}\kappa_{\eta,\,l}(c_{j}c_{j+l}+c_{j+l}^{\dagger}c_{j}^{\dagger}),$ (52)
@@
Clearly, QFI scales in a super-Heisenberg manner if the long distance couplings decay slower than Coulombic, i.e. $\eta\leq 1$.
@@ V.1 Pseudo-random localization transition
$\hat{H}$ $=$ $-t\sum_{i}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+h.c)+V\sum_{i}\cos(2\pi i\omega)\hat{c}^{\dagger}_{i}\hat{c}_{i}$ (54)
@@
where $t$ is tunneling, $\omega$ is an irrational number, $U$ is interaction strength, and $V$ is the unknown parameter which we wish to estimate. In addition, the operator $\hat{c}^{\dagger}_{i}(\hat{c}_{i})$ is fermionic creation (annihilation) operator at site $i$ and $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}_{i}$ accounts for particle number. The non-interacting case, i.e. $U=0$, is known as the Aubry-André-Harper (AAH) model Hofstadter and Mcintyre (1949). In the AAH model with a single particle, there is an energy independent localization transition at a finite modulation strength, $V_{c}=2$. For a given $V$, all the states are either localized (for $V>V_{c}$) or extended (for $V<V_{c}$ofstadter (1976). For finite size systems, proper scaling emerges at the transition for system sizes $F_{n}$ with either odd or even sequences from the Fibonacci series and for $\omega$ to be approximated by $\omega_{n}{=}F_{n}/F_{n+1}$. Here $F_{n}$ and $F_{n+1}$ are two consecutive Fibonacci numbers with the property $\omega{=}\lim_{n\to\infty}(F_{n}/F_{n+1})\to(\sqrt{5}-1)/2$, which is the so-called golden ratio.
@@ V.2 Stark localization
Stark localization caused by an applied gradient field across a lattice has been exploited for ultra-precise sensing He et al. (2023); Yousefjani et al. (2023a, 2024a); Sahoo and Rakshit (2024). The extended-localized transition in the limit of large one-dimensional systems takes place at an infinitesimal gradient field. This allows for sensing weak fields with great precision. This section is devoted to Stark probes and their capability to serve as a weak-gradient field sensor.
@@
$H=J\sum_{i=1}^{L-1}\left(|i\rangle\langle i+1|+|i+1\rangle\langle i|\right)+h\sum_{i=1}^{L}i|i\rangle\langle i|,$ (56)
@@
where $J$ is the tunneling rate. In this setup, the gradient-field strength $h$ is the unknown parameter and aimed to be estimated. It is well known that in the thermodynamic limit (i.e. $L{\rightarrow}\infty$), all energy levels of the Hamiltonian Eq.(56) go through a phase transition from an extended to a localized phase at $h_{c}{=}0$ van Nieuwenburg et al. (2019); Kolovsky (2008); Schulz et al. (2019); Chanda et al. (2020); Yao and Zakrzewski (2020). Focusing on the ground state, Fig. 5 (a) presents QFI as a function of $h$ for different sizes of the probe. The initial plateaus represent the extended phase in which the QFI takes its maximum for $h{\leq}h_{\max}$. Clearly, $h_{\max}$ skews to zero and tends to the critical point $h_{c}{=}0$, as the length of the probe increases. The maximum values of the QFI, denoted by $I_{Q}(h_{\max})$, increase dramatically by enlarging the probe, evidencing divergence of the QFI in the thermodynamic limit. In the localized phase, namely $h{>}h_{\max}$, the QFI is size-independent and decays algebraically according to $I_{Q}{\propto}|h-h_{\max}|^{-\alpha}$ with $\alpha{=}2$ for the ground state. The scaling of the QFI with the system size at the critical point has been plotted in Fig. 5 (b). The QFI shows strong super-Heisenberg scaling, i.e. $I_{Q}(h_{\max}){\propto}L^{\beta}$ with $\beta{\simeq}6$. The inset of Fig. 5 (b) shows how $\beta$ changes by getting distance from the critical point, i.e. $h_{\max}$. As the Stark localization transition happens in the entire energy spectrum of Eq. (56), one can inquire about higher energy eigenstates’ scaling. Interestingly, one can obtain the super-Heisenberg scaling with $\beta{\simeq}4$ for all excited eigenstates. Following section IV, establishing finite-size scaling analysis results in critical exponents $(\alpha,\nu){=}(2.00,0.33)$ and $(\alpha,\nu){=}(4.00,1.00)$ for the ground state and a typical eigenstate from the mid-spectrum, respectively. Indeed, three exponents $\beta,\alpha$ and $\nu$ obey $\beta{=}\alpha/\nu$ in Eq. (42), regardless of the energy level.
@@
Figure 5: Single-particle Stark probe; (a) QFI as a function of the Stark filed $h$ for single-particle Stark probe, prepared in the ground state of Eq. (56) with various sizes. The dashed line is the best fitting function to describe the algebraic behavior of the QFI in the localized regime when $L{\rightarrow}\infty$. (b) the maximum of the QFI at $h_{\max}$ versus probe size $L$. The numerical results (markers) are well-fitted by function $I_{Q}{\propto}L^{\beta}$ (line) with $\beta{=}5.98$. The inset shows the scaling behavior far from the criticality. Figure is adopted from He et al. (2023).
@@
Figure 6: Many-body Stark probe; (a) QFI versus $h$ for many-body Stark probe with nearest-neighbor interaction prepared in the ground state of Eq. (57) with $\eta{\rightarrow}\infty$ and different sizes. The dashed line which is the best-fitting function shows the asymptotic behavior of the QFI in the localized phase. (b) the QFI versus $L$ in extended phase, i.e. $h{=}10^{-4}J$, and critical point, i.e. $h{=}h_{\max}$. The numerical results (markers) are well-described by fitting function $I_{Q}{\propto}L^{\beta}$ (line) with $\beta{=}3.67$ and $\beta{=}4.26$ for the extended phase and critical point $h_{\max}$, respectively. (c) QFI versus $h$ for many-body Stark probe prepared in the ground state of Eq. (57) with different $\eta$ and $L{=}30$. (d) The obtained scales of the QFI, namely $\beta(h,\eta)$ versus range of the interaction $\eta$, in both extended phase ($h{=}10^{-4}$) and critical points ($h{=}h_{\max}$). Figure is adopted from He et al. (2023); Yousefjani et al. (2023a).
@@
The offered quantum-enhanced sensitivity by Stark localization is not limited to the non-interacting probe. To study the effect of the interaction, one can consider a one-dimensional probe of size $L$ containing $N{=}L{/}2$ particles with tunneling to their next neighbor and interaction range that decay algebraically depends on the exponent $\eta{>}0$. The Hamiltonian reads
$H=J\sum_{i=1}^{L-1}(\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1})+\sum_{i<j}\frac{J}{|i-j|^{\eta}}\sigma^{z}_{i}\sigma^{z}_{j}+h\sum_{i=1}^{L}i\sigma_{i}^{z},$ (57)
@@
with $J$ and $h$ as the coupling strength and the magnitude of the gradient filed, respectively. By changing the exponent $\eta$, one can smoothly interpolate between a fully connected graph for $\eta{=}0$ and a standard nearest-neighbor one-dimensional chain for $\eta{\rightarrow}\infty$. In the following, after presenting the results for the probe with short-range interaction (i.e. $\eta{\rightarrow}\infty$), we investigate the impact of long-range interaction (i.e. $\eta{<}\infty$) to address this question that whether longer range of interaction can provide more benefits for sensing tasks.
@@
Short-range interaction.-- Fig. 6(a) represents the QFI as a function of $h$ for various $L$ obtained for the ground state  He et al. (2023). By increasing the Stark field, the system goes through a phase transition from the extended phase to the localized one. Clearly, the QFI peaks at some $h_{\max}$ which gradually moves to zero, indicating that the Stark localization transition in the many-body interacting probe takes place at an infinitesimal gradient field. While finite-size effects are evident in the extended phase ($h{<}h_{\max}$), in the localized phase ($h{>}h_{\max}$), the QFI is size-independent and decays algebraically as $I_{Q}{\propto}|h-h_{\max}|^{-\alpha}$ with $\alpha{=}4.00$. The scaling behavior of the Stark probe has been presented in Fig. 6(b) for both extended phase ($h{=}10^{-4}J$) and transition point ($h{=}h_{\max}$). The best fitting function to describe the numerical results (markers) obtained as $I_{Q}{\propto}L^{\beta}$ (lines) with $\beta{=}3.67$ and $\beta{=}4.26$, for extended phase and transition point, respectively. Obviously, similar to the single particle probe, the many-body interacting probe also results in strong super-Heisenberg precision.
@@
Long-range interaction.-- To explore the advantages offered by long-range interactions in the Stark probe, Ref. Yousefjani et al. (2023a) presents the QFI for a Stark probe of size $L{=}30$ prepared in the ground state of Hamiltonian Eq. (57) with different $\eta$’s. The results are shown in Fig. 6(c). Regardless of the range of the interaction, increasing $h$ results in transition from extended phase to localize one, highlighted by a peak at $h_{\max}(\eta)$. Various $\eta$’s leaves different imprints on the QFI. By decreasing the range of the interaction the induced Zeeman energy splitting on each site changes from a uniform to nonuniform (resembling disorder) which cause diminution in QFI. The effect of nonuniform energy splitting gradually vanishes as the interaction tends to short-range. By considering a fitting function of the form $I_{Q}\sim L^{\beta}$, one can estimate the exponent $\beta$ for various choices of $(h,\eta)$. The results are presented in Fig. 6(d) which clearly shows the possibility of super-Heisenberg precision over a wide range of interactions.
@@
Capturing the ultimate precision limit provided by the QFI, and equally saturating the Cramér-Rao bound in Eq. (7), demands an experimental-friendly measurement setup. In Ref. He et al. (2023) it has been shown that a simple position measurement described by local projectors $\{\Pi_{i}=|i\rangle\langle i|\}_{i=1}^{L}$ results in CFI closely matches with the QFI in single particle probe. For many-body interacting case, the best observable to closely catch the QFI is obtained as $\mathcal{O}{=}\Pi_{i=1}^{L}\sigma_{i}^{z}$ Yousefjani et al. (2023a, 2024a).

### v3

@@ IV.4 Quantum many-body probes with long-range interactions
Since local interactions are useful in quantum-enhanced sensing, it is natural to ask whether criticality in models with longer-range interactions are also beneficial for metrology tasks Lakkaraju et al. (2023). Unfortunately, exact results are rare for such systems and approximate methods, such as Density Matrix Renormalization Group (DMRG), are often unsatisfactory for generic Hamiltonians with long-range interactions because of an area law violation. However, QFI has indeed been studied in many systems showing interesting physics. Perhaps the most studied model is the Lipshin-Meshkov-Glick (LMG) Hamiltonian for N sites in 1D, which generalizes the anisotropic XY Hamiltonian by dropping the nearest-neighbour assumption. The Hamiltonian can be written as
$\hat{H}_{P}=\frac{J}{N}\sum_{i,j,i<j}\left(\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}+\gamma\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{y}\right)+h\sum_{i}\hat{\sigma}_{i}^{z},$ (50)
where $\gamma$ is the anisotropy parameter, $J$ is exchange coupling, and $h$ is the magnetic field. This model has been solved through Bethe ansatz (Pan and Draayer, 1999; Morita et al., 2006), or alternatively by a Holstein-Primakoff transformation (Dusuel and Vidal, 2004, 2005). Considering $J$ as the unit of interaction, this model has three estimable parameters, the magnetic field strength $h$, anisotropy $\gamma$, and the temperature $T$ when system is in thermal equilibrium. At zero temperature, this system shows quantum criticality at $h{=}h_{c}{=}J$. In Ref. (Kwok et al., 2008), it has been shown that the QFI with respect to $h$ diverges at criticality as $I_{Q}\sim(h-h_{c})^{-2}$ (namely $\alpha{=}2$) in the symmetric phase ($h/J\geq 1$) and $I_{Q}\sim(h_{c}-h)^{-1/2}$ (namely $\alpha{=}1/2$) in the broken phase ($0\leq h/J\leq 1$). Remarkably, these exponents for QFI do not depend on the anisotropy $\gamma$. While at the critical point $h_{c}$, one obtains $I_{Q}(h_{c})\sim N^{4/3}$, indicating quantum-enhanced sensitivity, two-site reduced density matrices results in $I_{Q}(h_{c})\sim N^{2/3}$ (Ma et al., 2008). In (Salvatori et al., 2014), the authors considered the estimation of $\gamma$ and $T$. In the thermodynamic limit near criticality, the leading contribution to QFI for both scales as $I_{Q}(h_{c})\sim T^{2}$ at both phases, thus making this system attractive for relatively high-temperature thermometry. At this point, one should note that this also reflects a long-standing challenge for low-temperature thermometry, i.e., when $T\rightarrow 0$, the QFI tends to vanish and the estimation error diverges. Other approaches like engineering non-Markovian interactions bosonic probes have been proposed to overcome this Zhang et al. (2022b) divergence, although we are not aware of any such attempt in this many-body spin-chain context. Moreover, QFI for estimating anisotropy $\gamma$ in the thermodynamic limit shows a divergence at $\gamma=1$ and scales differently in the ordered phase ($h>1$) compared to the symmetry-broken phase ($10\leq h<$ 1) following the formula
@@
Garbe et al (Garbe et al., 2022) have recently considered the influence of various driving protocols on QFI-based sensing in LMG and other fully connected systems. In LMG and other fully connected models, all interactions are equally strong, in contrast to models where all interactions except the nearest or next nearest are ignored. Thus, one may wonder what happens for models where long-range interactions are present, but the interaction strength decays following a power law (Fernández-Lorenzo et al., 2018). Such models become extremely relevant for ion-trap based platforms, where the decay exponent can be fixed by tuning the lasers Monroe et al. (2021). In Ref. (Yang et al., 2022a), the authors considered a long-range Kitaev chain, which is an extension of the $N$-site tight-binding model, and whose Hamiltonian is given by
$\hat{H}$ $=-\frac{J}{2}\sum_{j=1}^{N}(\hat{c}_{j}^{\dagger}\hat{c}_{j+1}+\hat{c}_{j+1}^{\dagger}\hat{c}_{j})-\mu\sum_{j=1}^{N}(\hat{c}_{j}^{\dagger}\hat{c}_{j}-\frac{1}{2})$
$+\frac{\Delta}{2}\sum_{j=1}^{N-1}\sum_{l=1}^{N-j}\kappa_{\eta,\,l}(\hat{c}_{j}\hat{c}_{j+l}+\hat{c}_{j+l}^{\dagger}\hat{c}_{j}^{\dagger}),$ (52)
@@
Clearly, QFI shows quantum-enhanced sensitivity if the long-distance couplings decay slower than Coulombic, i.e. $\eta\leq 1$.
@@ V.1 Pseudo-random localization transition
$\hat{H}$ $=$ $-J\sum_{i}(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+h.c)+V\sum_{i}\cos(2\pi i\omega)\hat{c}^{\dagger}_{i}\hat{c}_{i}$ (54)
@@
where $J$ is tunneling, $\omega$ is an irrational number, $U$ is interaction strength, and $V$ is the unknown parameter which we wish to estimate. In addition, the operator $\hat{c}^{\dagger}_{i}(\hat{c}_{i})$ is fermionic creation (annihilation) operator at site $i$ and $\hat{n}_{i}=\hat{c}^{\dagger}_{i}\hat{c}_{i}$ accounts for particle number. The non-interacting case, i.e. $U=0$, is known as the Aubry-André-Harper (AAH) model Hofstadter and Mcintyre (1949). In the AAH model with a single particle, there is an energy independent localization transition at a finite modulation strength, $V_{c}=2$. For a given $V$, all the states are either localized (for $V>V_{c}$) or extended (for $V<V_{c}$ofstadter (1976). For finite size systems, proper scaling emerges at the transition for system sizes $F_{n}$ with either odd or even sequences from the Fibonacci series and for $\omega$ to be approximated by $\omega_{n}{=}F_{n}/F_{n+1}$. Here $F_{n}$ and $F_{n+1}$ are two consecutive Fibonacci numbers with the property $\omega{=}\lim_{n\to\infty}(F_{n}/F_{n+1})\to(\sqrt{5}-1)/2$, which is the so-called golden ratio.
@@ V.2 Stark localization
Stark localization caused by an applied gradient field across a lattice has been exploited for ultra-precise sensing He et al. (2023); Yousefjani et al. (2023a, 2025a); Sahoo and Rakshit (2024). The extended-localized transition in the limit of large one-dimensional systems takes place at an infinitesimal gradient field. This allows for sensing weak fields with great precision. This section is devoted to Stark probes and their capability to serve as a weak-gradient field sensor.
@@
$\hat{H}=J\sum_{i=1}^{L-1}\left(|i\rangle\langle i+1|+|i+1\rangle\langle i|\right)+\theta\sum_{i=1}^{L}i|i\rangle\langle i|,$ (56)
@@
where $J$ is the tunneling rate. In this setup, the gradient-field strength $\theta$ is the unknown parameter and aimed to be estimated. It is well known that in the thermodynamic limit (i.e. $L{\rightarrow}\infty$), all energy levels of the Hamiltonian Eq.(56) go through a phase transition from an extended to a localized phase at $\theta_{c}{=}0$ van Nieuwenburg et al. (2019); Kolovsky (2008); Schulz et al. (2019); Chanda et al. (2020); Yao and Zakrzewski (2020). Focusing on the ground state, Fig. 5 (a) presents QFI as a function of $h$ for different sizes of the probe. The initial plateaus represent the extended phase in which the QFI takes its maximum for $\theta{\leq}\theta_{\max}$. Clearly, $h_{\max}$ skews to zero and tends to the critical point $\theta_{c}{=}0$, as the length of the probe increases. The maximum values of the QFI, denoted by $I_{Q}(\theta_{\max})$, increase dramatically by enlarging the probe, evidencing divergence of the QFI in the thermodynamic limit. In the localized phase, namely $\theta{>}\theta_{\max}$, the QFI is size-independent and decays algebraically according to $I_{Q}{\propto}|\theta-\theta_{\max}|^{-\alpha}$ with $\alpha{=}2$ for the ground state. The scaling of the QFI with the system size at the critical point has been plotted in Fig. 5 (b). The QFI shows strong quantum-enhanced sensitivity, i.e. $I_{Q}(\theta_{\max}){\propto}L^{\beta}$ with $\beta{\simeq}6$. The inset of Fig. 5 (b) shows how $\beta$ changes by getting distance from the critical point, i.e. $\theta_{\max}$. As the Stark localization transition happens in the entire energy spectrum of Eq. (56), one can inquire about higher energy eigenstates’ scaling. Interestingly, one can obtain the quantum-enhanced sensitivity with $\beta{\simeq}4$ for all excited eigenstates. Following section IV, establishing finite-size scaling analysis results in critical exponents $(\alpha,\nu){=}(2.00,0.33)$ and $(\alpha,\nu){=}(4.00,1.00)$ for the ground state and a typical eigenstate from the mid-spectrum, respectively. Indeed, three exponents $\beta,\alpha$ and $\nu$ obey $\beta{=}\alpha/\nu$ in Eq. (42), regardless of the energy level.
@@
Figure 5: Single-particle Stark probe; (a) QFI as a function of the Stark filed $\theta$ for single-particle Stark probe, prepared in the ground state of Eq. (56) with various sizes. The dashed line is the best fitting function to describe the algebraic behavior of the QFI in the localized regime when $L{\rightarrow}\infty$. (b) the maximum of the QFI at $\theta_{\max}$ versus probe size $L$. The numerical results (markers) are well-fitted by function $I_{Q}{\propto}L^{\beta}$ (line) with $\beta{=}5.98$. The inset shows the scaling behavior far from the criticality. Figure is adapted from He et al. (2023).
@@
Figure 6: Many-body Stark probe; (a) QFI versus $\theta$ for many-body Stark probe with nearest-neighbor interaction prepared in the ground state of Eq. (57) with $\eta{\rightarrow}\infty$ and different sizes. The dashed line which is the best-fitting function shows the asymptotic behavior of the QFI in the localized phase. (b) the QFI versus $L$ in extended phase, i.e. $\theta{=}10^{-4}J$, and critical point, i.e. $\theta{=}\theta_{\max}$. The numerical results (markers) are well-described by fitting function $I_{Q}{\propto}L^{\beta}$ (line) with $\beta{=}3.67$ and $\beta{=}4.26$ for the extended phase and critical point $\theta_{\max}$, respectively. (c) QFI versus $\theta$ for many-body Stark probe prepared in the ground state of Eq. (57) with different $\eta$ and $L{=}30$. (d) The obtained scales of the QFI, namely $\beta(\theta,\eta)$ versus range of the interaction $\eta$, in both extended phase ($\theta{=}10^{-4}$) and critical points ($\theta{=}\theta_{\max}$). Figure is adapted from He et al. (2023); Yousefjani et al. (2023a).
@@
The offered quantum-enhanced sensitivity by Stark localization is not limited to the non-interacting probe. To study the effect of the interaction, one can consider a one-dimensional probe of size $L$ containing $N_{\rm ex}{=}L{/}2$ excitations with tunneling to their next neighbor and interaction range that decay algebraically depends on the exponent $\eta{>}0$. The Hamiltonian reads
$\hat{H}=J\sum_{i=1}^{L-1}(\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}+\hat{\sigma}^{y}_{i}\hat{\sigma}^{y}_{i+1})+\sum_{i<j}\frac{J}{|i-j|^{\eta}}\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}+\theta\sum_{i=1}^{L}i\hat{\sigma}_{i}^{z},$ (57)
@@
with $J$ and $\theta$ as the coupling strength and the magnitude of the gradient field, respectively. By changing the exponent $\eta$, one can smoothly interpolate between a fully connected graph for $\eta{=}0$ and a conventional nearest-neighbor one-dimensional chain for $\eta{\rightarrow}\infty$. In the following, after presenting the results for the probe with short-range interaction (i.e. $\eta{\rightarrow}\infty$), we investigate the impact of long-range interaction (i.e. $\eta{<}\infty$) to address this question that whether longer range of interaction can provide more benefits for sensing tasks.
@@
Short-range interaction.-- Fig. 6(a) represents the QFI as a function of $\theta$ for various $L$ obtained for the ground state  He et al. (2023). By increasing the Stark field, the system goes through a phase transition from the extended phase to the localized one. Clearly, the QFI peaks at some $\theta_{\max}$ which gradually moves to zero, indicating that the Stark localization transition in the many-body interacting probe takes place at an infinitesimal gradient field. While finite-size effects are evident in the extended phase ($\theta{<}\theta_{\max}$), in the localized phase ($\theta{>}\theta_{\max}$), the QFI is size-independent and decays algebraically as $I_{Q}{\propto}|\theta-\theta_{\max}|^{-\alpha}$ with $\alpha{=}4.00$. The scaling behavior of the Stark probe has been presented in Fig. 6(b) for both extended phase ($\theta{=}10^{-4}J$) and transition point ($\theta{=}\theta_{\max}$). The best fitting function to describe the numerical results (markers) obtained as $I_{Q}{\propto}L^{\beta}$ (lines) with $\beta{=}3.67$ and $\beta{=}4.26$, for extended phase and transition point, respectively. Obviously, similar to the single particle probe, the many-body interacting probe also results in strong quantum-enhanced sensitivity.
@@
Long-range interaction.-- To explore the advantages offered by long-range interactions in the Stark probe, Ref. Yousefjani et al. (2023a) presents the QFI for a Stark probe of size $L{=}30$ prepared in the ground state of Hamiltonian Eq. (57) with different $\eta$’s. The results are shown in Fig. 6(c). Regardless of the range of the interaction, increasing $\theta$ results in transition from extended phase to localize one, highlighted by a peak at $\theta_{\max}(\eta)$. Various $\eta$’s leaves different imprints on the QFI. By decreasing the range of the interaction the induced Zeeman energy splitting on each site changes from a uniform to nonuniform (resembling disorder) which causes diminution in QFI. The effect of nonuniform energy splitting gradually vanishes as the interaction tends to short-range. By considering a fitting function of the form $I_{Q}\sim L^{\beta}$, one can estimate the exponent $\beta$ for various choices of $(\theta,\eta)$. The results are presented in Fig. 6(d) which clearly shows the possibility of quantum-enhanced sensitivity over a wide range of interactions.
@@
Capturing the ultimate precision limit provided by the QFI, and equally saturating the Cramér-Rao bound in Eq. (7), demands an experimental-friendly measurement setup. In Ref. He et al. (2023) it has been shown that a simple position measurement described by local projectors $\{\hat{\Pi}_{i}=|i\rangle\langle i|\}_{i=1}^{L}$ results in CFI closely matches with the QFI in single particle probe. For many-body interacting case, the best observable to closely catch the QFI is obtained as $\hat{\mathcal{O}}{=}\hat{\Pi}_{i=1}^{L}\hat{\sigma}_{i}^{z}$ Yousefjani et al. (2023a, 2025a).